# Scalable Chain of Thoughts via Elastic Reasoning

### v1

@@ Abstract
Large reasoning models (LRMs) have achieved remarkable progress on complex tasks by generating extended chains of thought (CoT). However, their uncontrolled output lengths pose significant challenges for real-world deployment, where inference-time budgets on tokens, latency, or compute are strictly constrained. We propose Elastic Reasoning, a novel framework for scalable chain of thoughts that explicitly separates reasoning into two phases—thinking and solution—with independently allocated budgets. At test time, Elastic Reasoning prioritize that completeness of solution segments, significantly improving reliability under tight resource constraints. To train models that are robust to truncated thinking, we introduce a lightweight budget-constrained rollout strategy, integrated into GRPO, which teaches the model to reason adaptively when the thinking process is cut short and generalizes effectively to unseen budget constraints without additional training. Empirical results on mathematical (AIME, MATH500) and programming (LiveCodeBench, Codeforces) benchmarks demonstrate that Elastic Reasoning performs robustly under strict budget constraints, while incurring significantly lower training cost than baseline methods. Remarkably, our approach also produces more concise and efficient reasoning even in unconstrained settings. The code will be made publicly available after completing the necessary review process.
@@ 2.2 Length Control in Large Language models
Controlling the generation length of an LLM directly affects both latency and monetary cost at inference time. Earlier approaches to length control are designed mainly for general text generation [13, 42]. Typical methods include (i) manipulating positional encodings to achieve exact sequence lengths [3], (ii) modifying training objectives to penalize deviations from length targets [13, 32], and (iii) fine-tuning on instructions that explicitly state the desired output length [42]. Although effective for tasks such as summarization or constrained writing, these techniques generally aim to verbosity or enforce maximum-length limits, and overlook the intricate, step-by-step reasoning processes required for many reasoning tasks. Recent works have begun to exploring efficiency in reasoning by encouraging shorter chains [14, 2]; however, they typically lack mechanisms for precise, user-defined length targets that align with explicit compute budgets. One notable attempt, budget forcing [24], enforces strict token caps by truncating or padding with special tokens. This can yield incomplete reasoning or unnatural, forced outputs, ultimately harming both accuracy and interpretability. Additionally, L1 [1] uses reinforcement learning to let models dynamically allocate inference compute based on constraints provided in the prompt. Our approach doesn’t need to include length instructions in the prompt. Instead, we truncate reasoning trajectories to meet a given budget and train the model under these constraints via reinforcement learning.
@@ 3.2.1 Budget-Constrained Inference
where $|y|$ denotes the number of generated tokens. However, naïvely truncating the output often results in incomplete or missing $y^{\text{solution}}$, leading to invalid or unusable predictions.
@@ 3.2.2 Separate Budgeting for Thinking and Solution
  • The model begins generating thinking segment within a <think> block.
@@ 3.2.3 Budget-Constrained Rollout
In our training setup, we fix the budget pair to $(t^{*},s^{*})=(1\text{K},1\text{K})$ for simplicity and efficiency. Surprisingly, we find that the learned policy generalizes well to a wide range of unseen budget configurations at test time, without requiring any additional fine-tuning. As shown in Figure 1, the E1-Math-1.5B model achieves substantial improvements and generalizes robustly across various generation budgets. This indicates that Elastic Reasoning encourages the model to internalize a flexible reasoning strategy that adapts to different resource constraints.
@@ 4.1 Models and Datasets
Our base models are DeepScaleR-1.5B-Preview [21] and DeepCoder-14B-Preview [20], which are fine-tuned from DeepSeekR1-Distill-Qwen-1.5B and 14B [4] through iterative context lengthening. For training data, we follow the same datasets used in [21, 20]. In the math domain, the training set consists of AIME (1984-2023), AMC, Omni-Math [9], and STILL [23]. For code training, we use TACO [16], SYNTHETIC-1 [22], and LiveCodeBench (2023/05/01-2024/07/31) [12]. For evaluation, we use AIME 2024, MATH500 [11], AMC, Olympiad-Bench [9], and Minerva Math [15] for mathematical reasoning. For code-related tasks, we evaluate on LiveCodeBench (2024/08/01-2025/02/01) [12], Codeforces, and HumanEval+ [18].
4.2 Mathematical Reasoning Results
@@ 4.3 Code Reasoning Results
Figure 5: Pass@1 accuracy on LiveCodeBench under varying reasoning budgets. Both of the models inference with separate budgeting strategy.
@@ 4.4.1 Which part is enhanced after training?
Table 2: Ablation of enhanced thinking and solution on DeepScaleR-1.5B-Preview and E1-Math-1.5. Budget is in format ‘thinking+solution’ (in thousands of tokens).
DeepScalaR-1.5B E1-Math-1.5B Pass@1 (%)
@@ 4.4.2 Ablation of Training Budget $t^{*}$
4.4.3 Thinking Solution Distribution
@@ 4.4.4 Iterative Training
Table 3: Pass@1(%) on AIME 2024 across different budget configurations in two iterations

### v2

@@ Abstract
Large reasoning models (LRMs) have achieved remarkable progress on complex tasks by generating extended chains of thought (CoT). However, their uncontrolled output lengths pose significant challenges for real-world deployment, where inference-time budgets on tokens, latency, or compute are strictly constrained. We propose Elastic Reasoning, a novel framework for scalable chain of thoughts that explicitly separates reasoning into two phases—thinking and solution—with independently allocated budgets. At test time, Elastic Reasoning prioritizes the completeness of solution segments, significantly improving reliability under tight resource constraints. To train models that are robust to truncated thinking, we introduce a lightweight budget-constrained rollout strategy, integrated into GRPO, which teaches the model to reason adaptively when the thinking process is cut short and generalizes effectively to unseen budget constraints without additional training. Empirical results on mathematical (AIME, MATH500) and programming (LiveCodeBench, Codeforces) benchmarks demonstrate that Elastic Reasoning performs robustly under strict budget constraints, while incurring significantly lower training cost than baseline methods. Remarkably, our approach also produces more concise and efficient reasoning even in unconstrained settings. Our code has been made available at https://github.com/SalesforceAIResearch/Elastic-Reasoning.
@@ 2.2 Length control in large language models
Controlling the generation length of an LLM directly affects both latency and monetary cost at inference time. Earlier approaches to length control are designed mainly for general text generation [13, 42]. Typical methods include (i) manipulating positional encodings to achieve exact sequence lengths [3], (ii) modifying training objectives to penalize deviations from length targets [13, 32], and (iii) fine-tuning on instructions that explicitly state the desired output length [42]. Although effective for tasks such as summarization or constrained writing, these techniques generally aim to verbosity or enforce maximum-length limits, and overlook the intricate, step-by-step reasoning processes required for many reasoning tasks. Recent works have begun to explore efficiency in reasoning by encouraging shorter chains [14, 2]; however, they typically lack mechanisms for precise, user-defined length targets that align with explicit compute budgets. One notable attempt, budget forcing [24], enforces strict token caps by truncating or padding with special tokens. This can yield incomplete reasoning or unnatural, forced outputs, ultimately harming both accuracy and interpretability. Additionally, L1 [1] uses reinforcement learning to let models dynamically allocate inference compute based on constraints provided in the prompt. Our approach does not need to include length instructions in the prompt. Instead, we truncate reasoning trajectories to meet a given budget and train the model under these constraints via reinforcement learning.
@@ 3.2.1 Budget-constrained inference
where $|y|$ denotes the number of generated tokens. However, naively truncating the output often results in incomplete or missing $y^{\text{solution}}$, leading to invalid or unusable predictions.
@@ 3.2.2 Separate budgeting for thinking and solution
  • The model begins generating within a <think> block.
@@ 3.2.3 Budget-constrained rollout
In our training setup, we fix the budget pair to $(t^{*},s^{*})=(1\text{K},1\text{K})$ for simplicity and efficiency. Surprisingly, we find that the learned policy generalizes well to a wide range of unseen budget configurations at test time, without requiring any additional fine-tuning. As shown in Figure 1, the E1-Math-1.5B model achieves substantial improvements while generalizing robustly across various generation budgets. This indicates that Elastic Reasoning encourages the model to internalize a flexible reasoning strategy that adapts to different resource constraints.
@@ 4.1 Models and datasets
Our base models are DeepScaleR-1.5B-Preview [21] and DeepCoder-14B-Preview [20], which are fine-tuned from DeepSeekR1-Distill-Qwen-1.5B and 14B [4] through iterative context lengthening. For training data, we follow the same datasets used in [21, 20]. In the math domain, the training set consists of AIME (1984-2023), AMC, Omni-Math [9], and STILL [23]. For code training, we use TACO [16], SYNTHETIC-1 [22], and LiveCodeBench (2023/05/01-2024/07/31) [12]. For evaluation, we use AIME 2024, MATH500 [11], AMC, Olympiad-Bench [9], and Minerva Math [15] for mathematical reasoning. For code-related tasks, we evaluate on LiveCodeBench (2024/08/01-2025/02/01) [12], Codeforces, and HumanEval+ [18]. More training details are in Appendix B.
4.2 Mathematical reasoning rsults
@@ 4.3 Code reasoning results
Figure 5: Pass@1 accuracy on LiveCodeBench under varying reasoning budgets. Both of the models inference with separate budgeting.
@@ 4.4.1 Which part is enhanced after training?
Table 2: Ablation of enhanced thinking and solution on DeepScaleR-1.5B-Preview and E1-Math-1.5. Budget is in format ‘thinking+solution‘ (in thousands of tokens).
DeepScaleR-1.5B E1-Math-1.5B Pass@1 (%)
@@ 4.4.2 Ablation of training budget $t^{*}$
Figure 6: Ablation study on training reasoning budget $t^{*}$. We compare four settings: $t^{*}\in\{0.5\text{K},\ 1\text{K},\ 2\text{K},\ 3\text{K}\}$, while keeping the solution budget fixed at $a^{*}=1\text{K}$.
@@
4.4.3 Token allocation between thinking and solution
@@ 4.4.4 Iterative training
Table 3: Pass@1(%) on AIME 2024 across different budget configurations in two iterations.
@@ 5 Conclusion
Appendix A Limitations and broader impacts
While Elastic Reasoning offers a practical and scalable solution for length-controllable inference, it introduces a few limitations. First, our method assumes that the reasoning process can be meaningfully segmented into distinct thinking and solution phases, which may not hold for tasks with highly interleaved reasoning and answering (e.g., some forms of dialogue or commonsense reasoning). Second, our evaluation focuses primarily on math and code reasoning tasks; further investigation is needed to assess the effectiveness of Elastic Reasoning in domains such as scientific question answering or multi-hop retrieval.
Elastic Reasoning aims to improve the deployability and controllability of large reasoning models by enabling efficient inference under strict compute constraints. This can have positive downstream effects, such as reducing energy consumption, lowering latency for interactive applications, and enabling access to reasoning-capable models in resource-limited environments. At the same time, the ability to tailor model outputs based on budget constraints introduces new challenges: models may be optimized to generate shorter outputs without transparent communication of omitted reasoning, potentially affecting interpretability or user trust. Additionally, as reasoning models are increasingly used in high-stakes domains (e.g., education, law, healthcare), the separation between “thinking” and “solution” phases may require careful design to ensure that important justifications are not lost during inference-time truncation. We encourage future work to explore human-centered evaluations of controllable reasoning, and to develop standards for safely deploying budget-aware models in real-world applications.
Appendix B Training details
For GRPO training, we adopt the same hyperparameters as those used in DeepScaleR-1.5B-Preview and DeepCode-14B-Preview. For E1-Math-1.5B, we use a learning rate of $1\times 10^{-6}$ and a batch size of 128. The maximum context length is set to 1K tokens for the thinking segment and 1K tokens for the solution segment. Training is performed for 200 steps using the VeRL [sheng2024hybridflow] framework. For E1-Code-14B, we use the same learning rate and batch size. The context length configuration mirrors that of the math model: 1K tokens for thinking and 1K tokens for solution. Training is conducted for only 30 steps.